OpenStudy (anonymous):

Express the complex number in trigonometric form. -4

OpenStudy (jdoe0001):

\(\large { -4\implies \begin{array}{ccllll} -4&+0i\\ a&b \end{array}\to \begin{array}{llll} a=-4\\ b=0 \end{array}\qquad r[cos(\theta)+isin(\theta)] \\ \quad \\ \quad \\ r=\sqrt{a^2+b^2}\qquad tan(\theta)=\cfrac{b}{a}\implies \theta=tan^{-1}\left(\cfrac{b}{a}\right) }\)